Grade 9Mechanics

A train moving at constant velocity of 54km moves east ward for the 30minutes then due to north with same speed for the 40minutes. What is average velocity for the train during this run

surendra surendra
7 Years agoGrade 9

2 Answers

Arun
ApprovedApproved Tutor Answer7 Years ago
distance covered due east= 54km/hr * 1/2 hr=27 km 
distance covered due north=54km/hr * 4/6hr=36 km
displacement is the hypotenuse between starting and initial point.
=√27² + 36² = 45km
total time= 70 mins or 7/6 hrs
avg velocity=total displaceent/time=45*6/7=38.4 km/hr
so ∴ avg velocity is 38.4 km/hr.
 
 
Jampana charitha reddy
5 Years ago
Distance covered due to east 54*30minutes=54*1/2 hours =27
Distance due to North 54*40minutes=54*2/3 hours =18 total distance=√27+18=45 now total time=1/2+2/3=5/6 average velocity = total distance covered /time =45*6/5=38.4